OpenStudy (anonymous):

factor the difference of the two squares 16-49x^2

OpenStudy (jdoe0001):

\(\bf 16-49x^2\qquad \textit{notice that}\quad 16=4^2\qquad 49=7^2\qquad thus\\ \quad \\ 16-49x^2\implies 4^2-7^2x^2\implies 4^2-(7x)^2\\ \quad \\ \textit{now recall the difference of squares}\qquad a^2-b^2 = (a-b)(a+b)\) and expand likewise

OpenStudy (anonymous):

and the answer would be?

OpenStudy (jdoe0001):

well, expand it